Assume that a sample is used to estimate a population mean μ . Find the 99.9% confidence interval for a sample of size 52 with a mean of 80.8 and a standard deviation of 17.2. Enter your answer as an open-interval (i.e., parentheses) accurate to one decimal place (because the sample statistics are reported accurate to one decimal place).

99.9% C.I. =
Answer #1

Solution :

Given that,

Point estimate = sample mean = \bar x = 80.8

sample standard deviation = s = 17.2

sample size = n = 52

Degrees of freedom = df = n - 1 = 51

At 99.9% confidence level the t is ,

\alpha = 1 - 999% = 1 - 0.999 = 0.001

\alpha / 2 = 0.001 / 2 = 0.0005

t\alpha /2,df = t0.0005,24 = 3.492

Margin of error = E = t\alpha/2,df * (s /\sqrtn)

= 3.492* (17.2 / \sqrt 52)

= 8.3

The 99% confidence interval estimate of the population mean is,

\bar x - E < \mu < \bar x + E

80.8 - 8.3 < \mu < 80.8 + 8.3

72.5 < \mu < 89.1

99.9% C.I. = (72.5 , 89.1)